The velocity of waves on a string is 92 m/s . If the fundamental frequency is 475 Hz, what is the length of the string
1
Expert's answer
2019-04-19T10:32:49-0400
"\lambda =\frac{v}{f}=\frac{92}{475}=0.1937m"

The length of the string:


"L=0.5\lambda=0.5(0.1937)=0.0969m=9.69cm."